Assume the following information for Larkspur Corp.

Accounts receivable (beginning balance) $160,000
Allowance for doubtful accounts (beginning balance) 11,400
Net credit sales 934,000
Collections 920,000
Write-offs of accounts receivable 6,500
Collections of accounts previously written off 2,300


Uncollectible accounts are expected to be 7% of the ending balance in accounts receivable.

Determine the ending balance in Accounts Receivable and the unadjusted balance in Allowance for Doubtful Accounts.

Ending balance in Accounts Receivable

$enter a dollar amount

Unadjusted balance in Allowance for Doubtful Accounts

$enter a dollar amount

Prepare the entry to record bad debt expense for the period. (Credit account titles are automatically indented when the amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Determine the ending (adjusted) balance in Allowance for Doubtful Accounts.

Ending (adjusted) balance in Allowance for Doubtful Accounts $enter the Ending (adjusted) balance in Allowance for Doubtful Accounts in dollars
Answer #1

1) Ending balance in Account receivable = 160000+934000-920000-6500 = 167500

Unadjusted balance of allowance for doubtful accounts = 11400-6500+2300 = 7200

2) Adjusting entry

Account title and explanation Debit Credit
Bad debt expense (167500*7%-7200) 4525
Allowance for doubtful accounts 4525

3) Ending adjusted balance in allowance for doubtful accounts = $11725